# Universal activated aging and weak ergodicity breaking in spin and structural glasses

### v4

@@ V Aging in spin glass random energy models
For the G-REM, the Gaussian correction in Eq. (3) is non-negligible as shown in Fig. 2. Consequently, we need to use the full expression $\Pi_{w}(t_{\rm w})\sim H\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ to fit the data (see Fig. 3c). Here we have set $\hat{x}=T/T_{\rm WEB}$, $T_{\rm WEB}=T_{\rm c}/(1-W)\approx 3.6\,T_{\rm c}$ (see Eq. 4), where $W=bT_{\rm c}/a\approx 0.722$ with $a=1.12$ and $b=0.96$ from Fig. 2 and $T_{\rm c}\approx 0.849$ from the literature [33]. The fitted $H$ and $\alpha$ are compared to the theory in Fig. 7(a,b). The coefficient $k$ depends on NN as expected by the theory $k\sim 1/N$ (see Fig. 3d).
@@
An interesting prediction from the GTM theory is that the WEB does not necessarily co-occur with the spin glass transition: for the E-REM, $T_{\rm WEB}=T_{\rm c}$, while for the G-REM, $T_{\rm WEB}\approx 3.6\,T_{\rm c}$. The factor 3.6 in the G-REM is significantly larger than one - thus the GTM scenario ($T_{\rm WEB}\approx 3.6\,T_{\rm c}$) can be unambiguously tested against the original BTM scenario ($T_{\rm WEB}=T_{\rm c}$[22]. Below we directly verify the GTM scenario. In Fig. 4 we plot the MC data of $\Pi_{w}(t_{\rm w})$ for both E-REM and G-REM, with $t_{\rm w}$ rescaled by $C=\frac{w\hat{x}}{1+\hat{x}}$ to collapse the short-time behavior $\Pi_{w}^{\rm 0}(t_{\rm w})=1-Ct_{\rm w}$ at different $T$ (note that $C$ is not a fitting parameter). If WEB occurs, $\Pi_{w}(\tilde{t}_{\rm w})$ should not vanish in the $\tilde{t}_{\rm w}\to\infty$ limit, where $\tilde{t}_{\rm w}\equiv C\,{t}_{\rm w}$ is the reduced waiting time. We define a time scale $\tilde{\tau}$ such that $\Pi_{w}(\tilde{t}_{\rm w}=\tilde{\tau})=\Pi_{\rm thr}$ and choose a threshold $\Pi_{\rm thr}=0.2$. At sufficiently high $T$ (in the ergodic phase), $\tilde{\tau}=\tilde{\tau}^{0}$ is a constant, where $\tilde{\tau}^{0}$ is defined by $\Pi_{w}^{0}(\tilde{t}_{\rm w}=\tilde{\tau}^{0})=1-\tilde{\tau}^{0}=\Pi_{\rm thr}$. As shown in Fig. 4, with an increasing $T$, the ratio $R=\tilde{\tau}/\tilde{\tau}^{0}$ approaches one around $T_{\rm c}$ in E-REM and around $T_{\rm WEB}\approx 3.6\,T_{\rm c}$ in G-REM. According to the GTM scenario, we expect $R\to\infty$ below $T_{\rm WEB}$ in the thermodynamic limit, because $\tilde{\tau}$ should diverge in the WEB phase. This expectation is confirmed by finite-size analyses at a fixed $T=1.17\,T_{\rm c}$ slightly above $T_{\rm c}$ in both models (see Fig. 4d-inset). In G-REM, $R$ increases with NN at this temperature; in contrast, in E-REM, $R\approx 1$ is nearly invariant. This is direct evidence of WEB above $T_{\rm c}$ in the G-REM.
@@
Table 1: Characteristic temperatures. The spin glass transition temperature $T_{\rm c}$ is obtained by thermodynamic calculations for G-REM [33] and E-REM [55]. The divergence temperature $T_{\rm 0}$ is estimated from the VFT fitting of the $\alpha$-relaxation time $\tau_{\rm eq}\sim\exp\left(\frac{C}{T-T_{\rm 0}}\right)$ for the WCA model [56]. The MCT temperature $T_{\rm MCT}$ is estimated from the MCT fitting, $\tau_{\rm eq}\sim(T-T_{\rm MCT})^{-\gamma}$ for the WCA model [57], and $D\sim(T-T_{\rm MCT})^{\gamma}$ of the diffusion constant $D$ for amorphous $\rm{SiO}_{2}$ [58]. The onset temperature $T_{\rm onset}$ in the WCA model corresponds to the crossover temperature where the Stokes-Einstein relation, $\tau_{\rm eq}\sim D^{-1}$, breaks down [59]. Using four different criteria, Ref. [60] reports $T_{\rm onset}=0.6-0.7$ for the WCA model. Data without references are obtained by the current study (see the main text and Appendix for details).
@@
VI Aging in structural glasses
@@
To further examine the universality of the GTM theory, we simulate two structural glass formers: the WCA model as an example of simple molecular systems, and the amorphous silica ($\rm{SiO}_{2}$) model as an example of network glasses (see Appendixes C and D for details). Because efficient sampling of energy barriers in structural glasses remains technically challenging [40, 61, 62], we will only compare the simulated aging function $\Pi_{w}(t_{\rm w})$ to theoretical predictions (see the data in Fig. 5 with a fixed $w=1/2$). Note that $\Pi_{w}(t_{\rm w})$ is related to $F_{w}(t_{\rm w})$ via $F_{w}(t_{\rm w})=f\Pi_{w}(t_{\rm w})$, where $f$ is the nonergodicity parameter defined by the large-time plateau value of the equilibrium incoherent scattering function $f=\lim_{t\to\infty}F_{\rm eq}(t)$. The nonergodicity parameter $f$ is analogous to the Edwards-Anderson order parameter $q_{\rm EA}$ in spin glasses, both of which are determined by the typical size of the glass basins.
@@
We first extract $H$ and $\alpha$ from the aging data, and compare them to theoretical predictions. As shown in Fig. 5(a,d), the high-$T$ simulation results can be fitted to the form, $F_{w}(t_{\rm w})=\tilde{H}\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$, from which $\tilde{H}$ and $\alpha$ are obtained. Note that this expression only works in the intermediate-$t_{\rm w}$ regime (see Eq. 5), and thus the small-$t_{\rm w}$ (microscopic dynamics) and large-$t_{\rm w}$ (thermalization) data points should be excluded from the fitting. At lower $T$, a dip in $F_{w}(t_{\rm w})$, which might be attributed to the so-called boson peak [63], appears after the short-time processes, making the above fitting form inapplicable. Nevertheless, at these temperatures, $\tilde{H}$ can still be estimated since the plateau is well-defined. Then the relation $H=\tilde{H}/f$ is used to estimate $H$ (see Appendix E for the data of $f$). The values of $H$ and $\alpha$ are plotted in Fig. 7(a,b), as functions of $\hat{x}=T/T_{\rm WEB}$ with a fixed $w=1/2$. With a proper choice of $T_{\rm WEB}$, these data collapse with the GTM theoretical curves Eq. (2) and (6). Note that $T_{\rm WEB}$ is obtained from fitting the data to the arcsin law Eq. (2). A further consistent check is illustrated in Fig. 1d for the WCA model: using the same $T_{\rm WEB}$ and $f$, the theoretical curve $\tilde{H}(w,\hat{x}=0.36)=fH(w,\hat{x}=0.36)$, with $H$ given by Eq. (2), agrees with the simulation data $F^{\rm ag}(w)$. Together, Fig. 1d and 7(a,b) suggest that the theory works for different $T$ and $w=t/t_{\rm w}$.
@@
The finite-size dependence of the coefficient $k$ in the logarithmic term $k\ln t_{\rm w}$ reveals very interesting properties of activated processes. According to Eq. (5), we expect $k\sim 1/N$ in mean-field models. This NN essentially appears in the Gaussian term in Eq. (3). In non-mean-field systems such as the WCA and amorphous silica models, NN should be replaced by the size $n$ of clusters involved in the activation events, because $n$ determines the true scale of the energy barrier. Such clusters are also called “mosaics” in the RFOT [44]. If the system size NN is smaller than the characteristic size $n$, then all particles participate in the activated aging, and $k(N)$ would depend on NN. On the other hand, if $N>n$, $k(N)$ would be independent of NN because $n$ imposes a cutoff of the relevant size. In other words, the finite-size data $k(N)$ can provide an estimate of the characteristic size $n$: with an increasing NN, the data of $k(N)$ should reach a constant at $N=n$. Furthermore, we can obtain a length scale $\xi_{\rm ag}\sim n^{1/d}$. Note that $\xi_{\rm ag}$ is a static length scale because it is time-independent. This length essentially reflects properties of the energy landscape, but remarkably, it is extracted from aging dynamics.
@@
Based on the above consideration, we focus on the logarithmic part of the $F_{w}(t_{\rm w})$ data, and perform fitting using the form $F_{w}(t_{\rm w})\sim-k\ln t_{\rm w}$ (see Fig. 5b,e) to obtain $k(N)$ (see Fig. 5c,f). In the WCA model, $k\sim N^{-\beta}$ with $\beta\approx 0.1-0.4$ when $N<n$ (see Appendix F), and $k\sim{\rm constant}$ when $N>n$. From these data, the characteristic size $n$ and length $\xi_{\rm ag}\equiv\frac{1}{2}(n/\rho)^{1/3}$ are determined. Note that there is no adjustable parameter in the definition of $\xi_{\rm ag}$ - the coefficient $1/2$ follows the convention that the static length is defined as a half of the box length. In the amorphous silica model, the plateau in $k(N)$ is not observed within the largest simulated system size. This difference is very interesting, because it highlights the distinct nature of the activation energy in the two models. In large WCA systems ($N>n$), the barrier energy $\Delta E[n(T)]$ is determined by the $T$-dependent $n(T)$ (note that $n(T)$ grows with a decreasing $T$), but not the system size NN. In contrast, in the amorphous silica model $\Delta E(N)$ depends only on NN, which suggests that $\Delta E$ is a $T$-independent constant for a given system size. Correspondingly, the equilibrium relaxation time behaves as $\tau_{\rm eq}\sim\exp[\Delta E(T)/T]$ and $\tau_{\rm eq}\sim\exp(\Delta E/T)$ in the two models - this is consistent with the well-known fact that WCA and amorphous silica belong to fragile and strong glass formers respectively [64] (see Fig. 12a).
VII Growth of the static length in the WCA model
@@
In Fig. 6, the characteristic length $\xi_{\rm ag}$ in the WCA model is compared to two other static lengths. The first one is the PTS length $\xi_{\rm PTS}$ that characterizes the influence of fixed boundaries on the ergodicity of the bulk particles in the cavity [41]. The $\xi_{\rm PTS}$ data plotted in Fig. 6 are obtained from Ref. [42] for the WCA model at the same density $\rho=1.2$. The second one is the length scale $\xi_{\rm Hessian}$ used to collapse the data of the smallest eigenvalue of the Hessian matrix at different $T$ [43], which matches $\xi_{\rm PTS}$ as noticed in Ref. [65]. The length $\xi_{\rm Hessian}$ has been measured for the Kob-Anderson binary Lennard-Jones (KA-LJ) model [43] but not yet for the WCA model. Ref. [42] shows that the PTS lengths, $\xi_{\rm PTS}(T)$, for the two models (WCA and KA-LJ) collapse if $T$ is rescaled by a factor of $4/3$, $\xi_{\rm PTS}^{\rm WCA}(T)=\xi_{\rm PTS}^{\rm KA-LJ}(4T/3)$. Based on this observation, in Fig. 6 we plot $\xi_{\rm Hessian}$ of the KA-LJ model from  [43] as a function of $3T/4$. Surprisingly, we find that not only $\xi_{\rm Hessian}$ and $\xi_{\rm PTS}$ collapse on the same curve as expected [65], the length $\xi_{\rm ag}$ also collapses. The growth range of each single length is limited within a factor of $2\sim 3$, due to different numerical challenges: to obtain $\xi_{\rm Hessian}$ and $\xi_{\rm PTS}$ one needs to equilibrate the system, which is impossible at a low $T$; to obtain $\xi_{\rm ag}$, one needs to fit the logarithmic decay of the aging function (see Fig. 5), which becomes difficult at a high $T$ due to rapid thermalization. However, if we combine the three lengths, the variation covers one decade.
@@
VIII Universal behavior and a unified phase diagram of aging

### v5

@@ Results
For the G-REM, the Gaussian correction in Eq. (3) is non-negligible as shown in Fig. 3. Consequently, we need to use the full expression $\Pi_{w}(t_{\rm w})\sim H\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ to fit the data (see Fig. 2c). Here we have set $\hat{x}=T/T_{\rm WEB}$, $T_{\rm WEB}=T_{\rm c}/(1-W)\approx 3.6\,T_{\rm c}$ (see Eq. 4), where $W=bT_{\rm c}/a\approx 0.722$ with $a=1.12$ and $b=0.96$ from Fig. 3 and $T_{\rm c}\approx 0.849$ from the literature [33]. The fitted $H$ and $\alpha$ are compared to the theory in Fig. 7(a,b). The coefficient $k$ depends on $N$ as expected by the theory $k\sim 1/N$ (see Fig. 2d).
@@
An interesting prediction from the GTM theory is that the WEB does not necessarily co-occur with the spin glass transition: for the E-REM, $T_{\rm WEB}=T_{\rm c}$, while for the G-REM, $T_{\rm WEB}\approx 3.6\,T_{\rm c}$. The factor 3.6 in the G-REM is substantially larger than one - thus the GTM scenario ($T_{\rm WEB}\approx 3.6\,T_{\rm c}$) can be unambiguously tested against the original BTM scenario ($T_{\rm WEB}=T_{\rm c}$[22]. Below we directly verify the GTM scenario. In Fig. 4 we plot the MC data of $\Pi_{w}(t_{\rm w})$ for both E-REM and G-REM, with $t_{\rm w}$ rescaled by $C=\frac{w\hat{x}}{1+\hat{x}}$ to collapse the short-time behavior $\Pi_{w}^{\rm 0}(t_{\rm w})=1-Ct_{\rm w}$ at different $T$ (note that $C$ is not a fitting parameter). If WEB occurs, $\Pi_{w}(\tilde{t}_{\rm w})$ should not vanish in the $\tilde{t}_{\rm w}\to\infty$ limit, where $\tilde{t}_{\rm w}\equiv C\,{t}_{\rm w}$ is the reduced waiting time. We define a time scale $\tilde{\tau}$ such that $\Pi_{w}(\tilde{t}_{\rm w}=\tilde{\tau})=\Pi_{\rm thr}$ and choose a threshold $\Pi_{\rm thr}=0.2$. At sufficiently high $T$ (in the ergodic phase), $\tilde{\tau}=\tilde{\tau}^{0}$ is a constant, where $\tilde{\tau}^{0}$ is defined by $\Pi_{w}^{0}(\tilde{t}_{\rm w}=\tilde{\tau}^{0})=1-\tilde{\tau}^{0}=\Pi_{\rm thr}$. As shown in Fig. 4, with an increasing $T$, the ratio $R=\tilde{\tau}/\tilde{\tau}^{0}$ approaches one around $T_{\rm c}$ in E-REM and around $T_{\rm WEB}\approx 3.6\,T_{\rm c}$ in G-REM. According to the GTM scenario, we expect $R\to\infty$ below $T_{\rm WEB}$ in the thermodynamic limit, because $\tilde{\tau}$ should diverge in the WEB phase. This expectation is confirmed by finite-size analyses at a fixed $T=1.17\,T_{\rm c}$ slightly above $T_{\rm c}$ in both models (see Fig. 4d-inset). In G-REM, $R$ increases with $N$ at this temperature; in contrast, in E-REM, $R\approx 1$ is nearly invariant. This is direct evidence of WEB above $T_{\rm c}$ in the G-REM.
@@
Table 1: Characteristic temperatures. The spin glass transition temperature $T_{\rm c}$ is obtained by thermodynamic calculations for G-REM [33] and E-REM [55]. The divergence temperature $T_{\rm 0}$ is estimated from the VFT fitting of the $\alpha$-relaxation time $\tau_{\rm eq}\sim\exp\left(\frac{C}{T-T_{\rm 0}}\right)$ for the WCA model [56]. The MCT temperature $T_{\rm MCT}$ is estimated from the MCT fitting, $\tau_{\rm eq}\sim(T-T_{\rm MCT})^{-\gamma}$ for the WCA model [57], and $D\sim(T-T_{\rm MCT})^{\gamma}$ of the diffusion constant $D$ for amorphous $\rm{SiO}_{2}$ [58]. The onset temperature $T_{\rm onset}$ in the WCA model corresponds to the crossover temperature where the Stokes-Einstein relation, $\tau_{\rm eq}\sim D^{-1}$, breaks down [59]. Using four different criteria, Ref. [60] reports $T_{\rm onset}=0.6-0.7$ for the WCA model. Data without references are obtained by the current study (see the main text and SM Sec. S4 for details).
@@
Aging in structural glasses
@@
To further examine the universality of the GTM theory, we simulate two structural glass formers: the WCA model as an example of simple molecular systems, and the amorphous silica ($\rm{SiO}_{2}$) model as an example of network glasses (see Materials and Methods B and C for details). Because efficient sampling of energy barriers in structural glasses remains technically challenging [40, 61, 62], we will only compare the simulated aging function $\Pi_{w}(t_{\rm w})$ to theoretical predictions (see the data in Fig. 5 with a fixed $w=1/2$). Note that $\Pi_{w}(t_{\rm w})$ is related to $F_{w}(t_{\rm w})$ via $F_{w}(t_{\rm w})=f\Pi_{w}(t_{\rm w})$, where $f$ is the nonergodicity parameter defined by the large-time plateau value of the equilibrium incoherent scattering function $f=\lim_{t\to\infty}F_{\rm eq}(t)$. The nonergodicity parameter $f$ is analogous to the Edwards-Anderson order parameter $q_{\rm EA}$ in spin glasses, both of which are determined by the typical size of the glass basins.
@@
We first extract $H$ and $\alpha$ from the aging data, and compare them to theoretical predictions. As shown in Fig. 5(a,d), the high-$T$ simulation results can be fitted to the form, $F_{w}(t_{\rm w})=\tilde{H}\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$, from which $\tilde{H}$ and $\alpha$ are obtained. Note that this expression only works in the intermediate-$t_{\rm w}$ regime (see Eq. 5), and thus the small-$t_{\rm w}$ (microscopic dynamics) and large-$t_{\rm w}$ (thermalization) data points should be excluded from the fitting. At lower $T$, a dip in $F_{w}(t_{\rm w})$, which might be attributed to the so-called boson peak [63], appears after the short-time processes, making the above fitting form inapplicable. Nevertheless, at these temperatures, $\tilde{H}$ can still be estimated since the plateau is well-defined. Then the relation $H=\tilde{H}/f$ is used to estimate $H$ (see SM Sec. S4 for the data of $f$). The values of $H$ and $\alpha$ are plotted in Fig. 7(a,b), as functions of $\hat{x}=T/T_{\rm WEB}$ with a fixed $w=1/2$. With a proper choice of $T_{\rm WEB}$, these data collapse with the GTM theoretical curves Eq. (2) and (6). Note that $T_{\rm WEB}$ is obtained from fitting the data to the arcsin law Eq. (2). A further consistent check is illustrated in Fig. 1d for the WCA model: using the same $T_{\rm WEB}$ and $f$, the theoretical curve $\tilde{H}(w,\hat{x}=0.36)=fH(w,\hat{x}=0.36)$, with $H$ given by Eq. (2), agrees with the simulation data $F^{\rm ag}(w)$. Together, Fig. 1d and 7(a,b) suggest that the theory works for different $T$ and $w=t/t_{\rm w}$.
@@
The finite-size dependence of the coefficient $k$ in the logarithmic term $k\ln t_{\rm w}$ reveals very interesting properties of activated processes. According to Eq. (5), we expect $k\sim 1/N$ in mean-field models. This $N$ essentially appears in the Gaussian term in Eq. (3). In non-mean-field systems such as the WCA and amorphous silica models, $N$ should be replaced by the size $n$ of clusters involved in the activation events, because $n$ determines the true scale of the energy barrier. Such clusters are also called “mosaics” in the RFOT [44]. If the system size $N$ is smaller than the characteristic size $n$, then all particles participate in the activated aging, and $k(N)$ would depend on $N$. On the other hand, if $N>n$, $k(N)$ would be independent of $N$ because $n$ imposes a cutoff of the relevant size. In other words, the finite-size data $k(N)$ can provide an estimate of the characteristic size $n$: with an increasing $N$, the data of $k(N)$ should reach a constant at $N=n$. Furthermore, we can obtain a length scale $\xi_{\rm ag}\sim n^{1/d}$. Note that $\xi_{\rm ag}$ is a static length scale because it is time-independent. This length essentially reflects properties of the energy landscape, but remarkably, it is extracted from aging dynamics.
@@
Based on the above consideration, we focus on the logarithmic part of the $F_{w}(t_{\rm w})$ data, and perform fitting using the form $F_{w}(t_{\rm w})\sim-k\ln t_{\rm w}$ (see Fig. 5b,e) to obtain $k(N)$ (see Fig. 5c,f). In the WCA model, $k\sim N^{-\beta}$ with $\beta\approx 0.1-0.4$ when $N<n$ (see SM Sec. S4), and $k\sim{\rm constant}$ when $N>n$. From these data, the characteristic size $n$ and length $\xi_{\rm ag}\equiv\frac{1}{2}(n/\rho)^{1/3}$ are determined. Note that there is no adjustable parameter in the definition of $\xi_{\rm ag}$ - the coefficient $1/2$ follows the convention that the static length is defined as a half of the box length. In the amorphous silica model, the plateau in $k(N)$ is not observed within the largest simulated system size. This difference is very interesting, because it highlights the distinct nature of the activation energy in the two models. In large WCA systems ($N>n$), the barrier energy $\Delta E[n(T)]$ is determined by the $T$-dependent $n(T)$ (note that $n(T)$ grows with a decreasing $T$), but not the system size $N$. In contrast, in the amorphous silica model $\Delta E(N)$ depends only on $N$, which suggests that $\Delta E$ is a $T$-independent constant for a given system size. Correspondingly, the equilibrium relaxation time behaves as $\tau_{\rm eq}\sim\exp[\Delta E(T)/T]$ and $\tau_{\rm eq}\sim\exp(\Delta E/T)$ in the two models - this is consistent with the well-known fact that WCA and amorphous silica belong to fragile and strong glass formers respectively [64].
Growth of the static length in the WCA model
@@
In Fig. 6, the characteristic length $\xi_{\rm ag}$ in the WCA model is compared to two other static lengths. (i) The first one is the PTS length, $\xi_{\rm PTS}$, a standard measure of the mosaic size, which scales as $n\sim(\xi_{\rm PTS})^{d}$ [41]. In practice, starting from an equilibrium configuration, particles outside a cavity of radius $R$ are pinned. The characteristic decay length $\xi_{\rm PTS}$ is then extracted by measuring how the overlap between the initial and new equilibrium configurations decays as a function of $R$ [65, 66, 67]. Thus, $\xi_{\rm PTS}$ quantifies the range of influence exerted by the amorphous boundary. The $\xi_{\rm PTS}$ data plotted in Fig. 6 are obtained from Ref. [42] for the WCA model at the same density $\rho=1.2$. (ii) The second one is the length scale $\xi_{\rm Hessian}$ used to collapse the data of the smallest eigenvalue of the Hessian matrix at different $T$ [43], which matches $\xi_{\rm PTS}$ as noticed in Ref. [68]. The length $\xi_{\rm Hessian}$ has been measured for the Kob-Anderson binary Lennard-Jones (KA-LJ) model [43] but not yet for the WCA model. Ref. [42] shows that the PTS lengths, $\xi_{\rm PTS}(T)$, for the two models (WCA and KA-LJ) collapse if $T$ is rescaled by a factor of $4/3$, $\xi_{\rm PTS}^{\rm WCA}(T)=\xi_{\rm PTS}^{\rm KA-LJ}(4T/3)$. Based on this observation, in Fig. 6 we plot $\xi_{\rm Hessian}$ of the KA-LJ model from  [43] as a function of $3T/4$.
Surprisingly, we find that not only $\xi_{\rm Hessian}$ and $\xi_{\rm PTS}$ collapse on the same curve as expected [68], the length $\xi_{\rm ag}$ also collapses. The growth range of each single length is limited within a factor of $2\sim 3$, due to different numerical challenges: to obtain $\xi_{\rm Hessian}$ and $\xi_{\rm PTS}$ one needs to equilibrate the system, which is impossible at a low $T$; to obtain $\xi_{\rm ag}$, one needs to fit the logarithmic decay of the aging function (see Fig. 5), which becomes difficult at a high $T$ due to rapid thermalization. However, if we combine the three lengths, the variation covers one decade. We have examined that crystallization and fractionation are absent within the aging dynamics under investigation, and therefore confirmed that the growth of $\xi_{\rm ag}$ is not due to these effects (see SM Sec. S4.)
@@
Universal behavior and a unified phase diagram of aging